# Complexity of Finite Borel Asymptotic Dimension

### v1

@@ 5.2. From hitting sets to finite asymptotic dimension
for every $x\in T_{z}$, where $0\leq k\leq 2m$ is the unique integer such that $k(x)\in I_{k}$. Note that $c_{H}:X\to\{0,1\}$ is Borel.
@@
Let $t\in\mathbb{N}^{+}$, $f:X\to X$ be an acyclic Borel function and $H$ be a Borel $4(6t)^{2}$-forward independent hitting set. Define $c^{-1}_{H}(\{i\})=U_{i}$ for each $i\in\{0,1\}$, where $c_{H}$ is defined as in the previous paragraph. Then $U_{0}$ and $U_{1}$ are Borel sets such that $X=U_{0}\sqcup U_{1}$ and for each $i\in\{0,1\}$, every equivalence class of $\mathcal{F}_{t}(U_{i})$ has diameter bounded by $728t+$ 7.
@@
Define $\ell(x)\in\mathbb{N}^{+}$ to be minimal such that $c_{H}(x)\not=c_{H}(f^{\ell(x)}(x))$ and set
@@
As $e(x)=e(f^{\ell(x)-1}(x))$, we may assume without loss of generality that $\ell(x)=1$. First, we claim that $f^{j}(x)\in U_{1-i}\setminus H$ for every $11\leq j\leq\frac{s}{3}+$ 1.
Indeed, if $f^{j}(x)\in H$ for any $11\leq j\leq\frac{s}{3}+$ 1, then, by the definition of $c_{H}$, we have $x\in U_{1-i}$, a contradiction. Similarly, assume for contradiction that $11\leq j\leq\frac{s}{3}+$ 1 is minimal such that $f^{j}(x)\not\in U_{1-i}$. Then $j\geq 2$, $f^{j-1}(x)\in U_{1-i}$ and we have that $x\in U_{1-i}$ as either $\lfloor k(x)/s\rfloor=\lfloor k(f^{j-1}(x))/s\rfloor\mod 2$ or $k(x),k(f^{j-1}(x))\in I_{k}$ for some $0\leq k\leq 2\ell$, where $I_{k}$ is from (5.1).
We finish the proof by showing that for every $0\leq j\leq\frac{s}{3}$ and $z\in f^{-j}(e(x))$, we have that $z\in U_{1-i}$. This clearly holds when $k(z)=k(f^{\frac{s}{3}+1-j}(x))$ by the previous paragraph. In particular, as $H$ is $r$-forward independent, it holds whenever $z\in H$. Now, the claim follows as, by the definition of $c_{H}$, we have that $c_{H}(w)=c_{H}(z)$ whenever $f^{m}(w)=z\in H$ and $10\leq m\leq s-$ 1. ∎
@@
Let $t\in\mathbb{N}^{+}$, $f:X\to X$ be an acyclic Borel function and $H$ be a Borel $4(6t)^{2}$-forward independent hitting set. Then there is a Borel equivalence relation $E$ on $X$ such that the $\rho_{f}$-diameter of every $E$-class is bounded by $728t+$ 7 and $B_{t}(x)$ meets at most $2$ $E$-classes for any $x\in X$.
@@
For each $x\in X$, set
@@
$E=(L\circ f^{t},L\circ f^{t})^{-1}(\text{Diag}_{X}).$
@@
$[x]_{E}\subseteq\bigcup_{j\leq t+(2s+2)}f^{-j}(L(f^{t}(x))).$
Indeed, if $y\in[x]_{E}$, then $L(f^{t}(y))=L(f^{t}(x))$. So the distance from $y$ to $L(f^{t}(x))$ is at most $t+(2s+2)$. Therefore, the diameter of $[x]_{E}$ is at most $2(t+(2s+2))$.
We finish the proof by showing that $B_{t}(x)$ meets at most $2$ $E$-classes for each $x\in X$. Note that $f^{t}[B_{t}(x)]=\{x,f(x),\dots,f^{2t}(x)\}$ and that
@@
(i’) $\implies$ (iii). Suppose that $\text{asdim}^{\text{eq}}_{B}(\mathcal{G}_{f})=d\in\mathbb{N}$. Let $t\in\mathbb{N}^{+}$, let $r=2t(d+1)$, and let $E$ be as in Definition 5.1. Then by Proposition 5.3, there is a Borel $t$-forward-independent hitting set for $f$, which gives (iii).
@@
(iii) $\implies$ (ii’). Given $t\in\mathbb{N}^{+}$, apply Proposition 5.8 to obtain a uniformly $\rho_{f}$-bounded Borel equivalence relation $E$ on $X$ such that $B_{t}(x)$ meets at most $2$ $E$-classes for any $x\in X$. ∎
@@
It is easy to show that $C$ is Borel and $f$-invariant and that the connectedness relation of $\mathcal{G}_{f}\restriction C$ admits a Borel transversal $T$ that is hitting for $f\restriction C$. From $T$, one can construct a witness to $\text{asdim}_{B}(\mathcal{G}_{f}\restriction\bigcup C)\leq 1$. Note, however, that for any $r\in\mathbb{N}^{+}$, if there is an $r$-forward-independent set for $f$, then $r\leq\ell$, where $\ell$ is the minimal length of a cycle in $\mathcal{G}_{f}$.
@@ 6.2. Hypersmoothness of graphs generated by Borel functions
The graph generated by a Borel function on a standard Borel space is hypersmooth (see [DJK94, Theorem 8.1]). The following is a well-known open problem.
@@ 6.3. A characterization of the complexity of $\operatorname{CSP}_{B}(H)$
Let $H$ be a sinkless and sourceless digraph. In Remark 1.5, we demonstrate that if $\operatorname{CSP}_{B}^{\operatorname{function}}(H^{p})$ is $\mathbf{\Sigma}^{1}_{2}$-complete for some abstract walk $p$, then $\operatorname{CSP}_{B}(H)$ is also $\mathbf{\Sigma}^{1}_{2}$-complete. Although there is already a complete algebraic understanding of when $\operatorname{CSP}_{B}(H)$ is $\mathbf{\Sigma}^{1}_{2}$-complete [BKN09, Tho22], a positive answer to the following problem, in combination with the proof of Theorem 1.4, would give a new perspective on this result.
@@ Acknowledgments
We would like to thank Zoltán Vidnyánszky, Alex Kastner, Andrew Marks, Sumun Iyer, and Forte Shinko for helpful discussions. In particular, we thank Alex Kastner for providing his note [Kas24] and a proof of Remark 3.4 and Andrew Marks for helping to simplify the proof of Theorem 3.1.
JG was supported by MSCA Postdoctoral Fellowships 2022 HORIZON-MSCA-2022-PF-01-01 project BORCA grant agreement number 101105722.

### v2

@@ 5.2. From hitting sets to finite asymptotic dimension
for every $x\in T_{z}$, where $0\leq k\leq 2m$ is the unique integer such that $k(x)\in I_{k}$. Note that $c_{H}:X\to\{0,1\}$ is Borel and that $\{x:c_{H}(x)\neq c_{H}(f(x))\}$ is a Borel (s1)(s-1)-forward-independent set.
@@
Let $t\in\mathbb{N}^{+}$, $f:X\to X$ be an acyclic Borel function and $H$ be a Borel $4(6t)^{2}$-forward independent hitting set. Define $c^{-1}_{H}(\{i\})=U_{i}$ for each $i\in\{0,1\}$, where $c_{H}$ is defined as in the previous paragraph. Then $U_{0}$ and $U_{1}$ are Borel sets such that $X=U_{0}\sqcup U_{1}$ and for each $i\in\{0,1\}$, every equivalence class of $\mathcal{F}_{t}(U_{i})$ has diameter bounded by $28t+7$.
@@
Recall from the definition of $c_{H}$ that $s=6t$. Define $\ell(x)\in\mathbb{N}^{+}$ to be minimal such that $c_{H}(x)\not=c_{H}(f^{\ell(x)}(x))$ and set
@@
As $e(x)=e(f^{\ell(x)-1}(x))$, we may assume without loss of generality that $\ell(x)=1$. First, we claim that $f^{j}(x)\in U_{1-i}\setminus H$ for every $1\leq j\leq\frac{s}{3}+1$.
Indeed, if $f^{j}(x)\in H$ for any $1\leq j\leq\frac{s}{3}+1$, then, by the definition of $c_{H}$, we have $x\in U_{1-i}$, a contradiction. Similarly, assume for contradiction that $1\leq j\leq\frac{s}{3}+1$ is minimal such that $f^{j}(x)\not\in U_{1-i}$. Then $j\geq 2$, $f^{j-1}(x)\in U_{1-i}$ and we have that $x\in U_{1-i}$ as either $\lfloor k(x)/s\rfloor=\lfloor k(f^{j-1}(x))/s\rfloor\mod 2$ or $k(x),k(f^{j-1}(x))\in I_{k}$ for some $0\leq k\leq 2\ell$, where $I_{k}$ is from (5.5).
We finish the proof by showing that for every $0\leq j\leq\frac{s}{3}$ and $z\in f^{-j}(e(x))$, we have that $z\in U_{1-i}$. This clearly holds when $k(z)=k(f^{\frac{s}{3}+1-j}(x))$ by the previous paragraph. In particular, as $H$ is $r$-forward independent, it holds whenever $z\in H$. Now, the claim follows as, by the definition of $c_{H}$, we have that $c_{H}(w)=c_{H}(z)$ whenever $f^{m}(w)=z\in H$ and $0\leq m\leq s-1$. ∎
@@
Let $t\in\mathbb{N}^{+}$, $f:X\to X$ be an acyclic Borel function and $H$ be a Borel $4(6t)^{2}$-forward independent hitting set. Then there is a Borel equivalence relation $E$ on $X$ such that the $\rho_{f}$-diameter of every $E$-class is bounded by $28t+7$ and $B_{t}(x)$ meets at most 22 $E$-classes for any $x\in X$.
@@
We continue using the notation developed in the definition of $c_{H}$ and the proof of Proposition 5.6. For each $x\in X$, set
@@
$E=(L\circ f^{t},L\circ f^{t})^{-1}(\text{Diag}_{X\times X}).$
@@
$x]_{E}\subseteq\bigcup_{j\leq t+(2s+2)}f^{-j}(L(f^{t}(x))).$
Indeed, if $y\in[x]_{E}$, then $L(f^{t}(y))=L(f^{t}(x))$. So the distance from $y$ to $L(f^{t}(x))$ is at most $t+(2s+2)$. Therefore, the diameter of $x]_{E}$ is at most $2(t+(2s+2))$.
We finish the proof by showing that $B_{t}(x)$ meets at most 22 $E$-classes for each $x\in X$. Note that $f^{t}[B_{t}(x)]=\{x,f(x),\dots,f^{2t}(x)\}$ and that
@@
(i’) $\implies$ (iii). Suppose that $\text{asdim}^{\text{eq}}_{B}(\mathcal{G}_{f})=d\in\mathbb{N}$. Let $r\in\mathbb{N}^{+}$ and set $r^{\prime}=2r(d+1)$. By Definition 5.1, there is a uniformly $\rho_{f}$-bounded Borel equivalence relation $E$ on $X$ such that, for each $x\in X$, $B_{r^{\prime}}(x)$ meets at most $(d+1)$ $E$-classes. Then by Proposition 5.3, applied with $t=r$, there is a Borel $r$-forward-independent hitting set for $f$, which gives (iii).
@@
(iii) $\implies$ (ii’). Given $t\in\mathbb{N}^{+}$, apply Proposition 5.9 to obtain a uniformly $\rho_{f}$-bounded Borel equivalence relation $E$ on $X$ such that $B_{t}(x)$ meets at most 22 $E$-classes for any $x\in X$. ∎
Remark 5.11.
One of the anonymous referees pointed out that there is another interesting equivalent condition that can be added to Theorem 5.10, which simplifies the proof of (iii) $\implies$ (ii’):
  1. (iv)
    For each $r>0$, there is a Borel map $g:X\to X$ such that $\sup_{x\in X}\rho_{f}(x,g(x))<\infty$ and $|g(B_{r}(x))|\leq 2$ for all $x\in X$.
We include a proof of the equivalence for completeness.
Proof.
(iii) $\implies$ (iv). Let $r\in\mathbb{N}$. We put $s=2r$ and assume without loss of generality that $s=6t$ for some $t>0$. Define $r^{\prime}=4s^{2}$. By (iii), there is a Borel $r^{\prime}$-forward-independent hitting set $H$. Define $c_{H}$ as in Subsection 5.2, and let $S=\{x\in X:c_{H}(x)\neq c_{H}(f(x))\}$. As noted before Proposition 5.6, $S$ is a Borel $s$-forward-independent set. For each $x\in X$, define $j(x)\in\mathbb{N}$ to be minimal such that $f^{j(x)}(x)\in S$. The argument in the first paragraph of the proof of Proposition 5.6 shows that $0\leq j(x)\leq 2s+1$ for every $x\in X$. Set $h(x)=f^{j(x)}(x)$, and observe that $h:X\to X$ is a Borel function. Define $g(x)=h(f^{s}(x))$. Clearly $g$ is Borel and $\sup_{x\in X}\rho_{f}(x,g(x))<\infty$ as we have $\rho_{f}(x,g(x))\leq s+(2s+1)=3s+1$ for every $x\in X$. Given $x\in X$, note that $g(B_{r}(x))=h(\{x,f(x),\dots,f^{s/2}(x)\})$, which has cardinality at most $2$ since $S$ is (s1)(s-1)-forward-independent and s1>s/2s-1>s/2. Then $g$ is as desired.
(iv) $\implies$ (ii’). Let $r\in\mathbb{N}$ and take $g$ as in (iv) with parameter $r\in\mathbb{N}$. Define an equivalence relation $E$ on $X$ by $(x,y)\in E$ if and only if $g(x)=g(y)$. As $E=(g\times g)^{-1}(\text{Diag}_{X\times X})$, we have that $E$ is Borel. To see that $E$ is $\rho_{f}$-uniformly bounded, observe that, for every $x\in X$, the diameter of $x]_{E}$ is at most $2M$, where $M$ is such that $\sup_{x\in X}\rho_{f}(x,g(x))=M$. Finally, since $g(B_{r}(x))$ has cardinality at most $2$ for every $x\in X$, we conclude that $B_{r}(x)$ meets at most 22 $E$-classes, and thus $\text{asdim}^{\text{eq}}_{B}(\mathcal{G}_{f})=1$. ∎
@@
It is easy to show that $C$ is Borel and $f$-closed and that the connectedness relation of $\mathcal{G}_{f}\restriction C$ admits a Borel transversal $T$ that is hitting for $f\restriction C$. From $T$, one can construct a witness to $\text{asdim}_{B}(\mathcal{G}_{f}\restriction C)\leq 1$. Note, however, that for any $r\in\mathbb{N}^{+}$, if there is an $r$-forward-independent set for $f$, then $r\leq\ell$, where $\ell=0$ if $f$ has a fixed point and $\ell$ is the minimal length of a cycle in $\mathcal{G}_{f}$ if $f$ has no fixed points.
@@ 6.2. Hypersmoothness of graphs generated by Borel functions
The graph generated by a Borel function on a standard Borel space is hypersmooth (see [13, Theorem 8.1]). The following is a well-known open problem.
@@ 6.3. A characterization of the complexity of $\operatorname{CSP}_{B}(H)$
Let $H$ be a sinkless and sourceless digraph. In Remark 1.5, we demonstrate that if $\operatorname{CSP}_{B}^{\operatorname{function}}(H^{p})$ is $\mathbf{\Sigma}^{1}_{2}$-complete for some abstract walk $p$, then $\operatorname{CSP}_{B}(H)$ is also $\mathbf{\Sigma}^{1}_{2}$-complete. Although there is already a complete algebraic understanding of when $\operatorname{CSP}_{B}(H)$ is $\mathbf{\Sigma}^{1}_{2}$-complete [1, 40], a positive answer to the following problem, in combination with the proof of Theorem 1.4, would give a new perspective on this result.
@@ Acknowledgments
We would like to thank Zoltán Vidnyánszky, Alex Kastner, Andrew Marks, Sumun Iyer, and Forte Shinko for helpful discussions. In particular, we thank Alex Kastner for providing his note on [41] and a proof of Remark 3.8 and Andrew Marks for helping to simplify the proof of Theorem 3.1. We thank the anonymous referees for helpful comments and suggestions. In particular, we thank one of the anonymous referees for pointing out an additional equivalent condition in Theorem 5.10, and for allowing us to include the suggested proof as Remark 5.11.
JG was supported by MSCA Postdoctoral Fellowships 2022 HORIZON-MSCA-2022-PF01-01 project BORCA grant agreement number 101105722, and by the Alexander von Humboldt Foundation in the framework of the Alexander von Humboldt Professorship of Daniel Kráľ endowed by the Federal Ministry of Education and Research.